OpenStudy (flashp):

Which of the following exponential functions goes through the points (1, 6) and (2, 12) f(x)=3(2)^x f(x)=2(3)^x f(x)=3(2)^-x f(x)=2(3)^-x

OpenStudy (dan815):

just plug in the point

OpenStudy (dan815):

plug in both points see which one goes through both

OpenStudy (flashp):

I did and I'm getting A, I just want to know if it's right or wrong

OpenStudy (flashp):

My answer is A

OpenStudy (dan815):

yes it is right

OpenStudy (waleed_imtiaz):

putting the point in all the options, I got A as correct

OpenStudy (flashp):

Thanks, someone who can finally help. One more question can you help me with it?

OpenStudy (dan815):

f(x)=3(2)^x 6=3*2^1, so 1,6 is a solution 12=3*2^2 so, 2*,12 is a solution

OpenStudy (flashp):

Two separate bacteria populations grow each month and are represented by the functions f(x) = 3x and g(x) = 7x + 6. In what month is the f(x) population greater than the g(x) population? Month 1 Month 2 Month 3 Month 4

OpenStudy (dan815):

should that say 3^x and 7^(x+6)

OpenStudy (dan815):

oh then its 7x+6 and 3^x?

OpenStudy (flashp):

As you see it is as it was wrote

OpenStudy (dan815):

impossible that is not your question

OpenStudy (flashp):

i swear it is

OpenStudy (dan815):

it was not copy pasted right go look at your question or take a screneshot for me

OpenStudy (dan815):

if that is your question there is simply no answer

OpenStudy (flashp):

That's it, word for word exactly how it was wrote

OpenStudy (dan815):



OpenStudy (dan815):

no its 3^x trust me

OpenStudy (dan815):

does it look like this \[3^x\]

OpenStudy (flashp):

whatever you think it is go for it, but it has to be either Month 1, Month 2, Month 3, or Month 4

OpenStudy (flashp):

Yeah it does look like that

OpenStudy (flashp):

Here's how it looks: f(x)=3^x and G(x)=7x+6 @dan815

OpenStudy (dan815):

take a screenshot of your question as it is

OpenStudy (dan815):

thats better

OpenStudy (dan815):